I'm not sure how to answer this question can anyone help me on this? Thanks

What is the extended delta update rule?
Answer #1

In machine learning, the Delta update rule is a gradient descent learning rule for updating the weights of the inputs to artificial neurons in a single-layer neural network. It is a special case of the more general backpropagation algorithm. For a neuron { j}j with activation function {g(x)}{\displaystyle g(x)}, the delta rule for {j}j's {i}ith weight {w_{ji}}{\displaystyle w_{ji}} is given by

{\displaystyle \Delta w_{ji}=\alpha (t_{j}-y_{j})g'(h_{j})x_{i}},

where

\alpha is a small constant called learning rate
{\displaystyle g(x)} is the neuron's activation function
{\displaystyle t_{j}} is the target output
{\displaystyle h_{j}} is the weighted sum of the neuron's inputs
{\displaystyle y_{j}} is the actual output
x_{i} is the {\displaystyle i}ith input.

It holds that {{j}=\sum x_{i}w_{ji}}{\displaystyle h_{j}=\sum x_{i}w_{ji}} and {y_{j}=g(h_{j})}{\displaystyle y_{j}=g(h_{j})}.

The delta rule is commonly stated in simplified form for a neuron with a linear activation function as

{\Delta w_{ji}=\alpha (t_{j}-y_{j})x_{i}}{\displaystyle \Delta w_{ji}=\alpha (t_{j}-y_{j})x_{i}}

While the delta rule is similar to the perceptron's update rule, the derivation is different. The perceptron uses the Heaviside step function as the activation function {\displaystyle g(h)}g(h), and that means that {\displaystyle g'(h)}g'(h) does not exist at zero, and is equal to zero elsewhere, which makes the direct application of the delta rule impossible.